OpenStudy (anonymous):

find solutions of the trigonometric equation in the interval [0,pi). 4cos^2 \[Phi]-1=0

OpenStudy (sweetnikkie072):

i would love to try to help you out..but i am unable to understand the equation. is it: \[4\cos ^{2}\div (\pi-1) or (4\cos ^{2}\div \pi)-1\] ?